OpenStudy (anonymous):

Convert the equation to the standard form for a parabola by completing the square on x or y as appropriate. y^2 + 2y - 2x - 3 = 0 A. (y + 1)^2 = 2(x + 2) B. (y - 1)^2 = -2(x + 2) C. (y + 1)^2 = 2(x - 2) D. (y - 1)^2 = 2(x + 2)

OpenStudy (sirm3d):

what is the "standard form"?

OpenStudy (anonymous):

y^2+2y+1-1 -2x-3 =0 (y+1)^2 -2x-4=0 (y+1)^2 = 2x+4 (y+1)^2 = 2(x+2)--------> option A is the right answer

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome